Suppose Q is the midpoint of segment PR. PQ = 2x + 1 and QR = 5x - 44. Find PQ.


1
Expert's answer
2020-08-26T17:39:32-0400

Given that Q is the midpoint of PR means,

PQ = QR


But, PQ = 2x + 1 and QR = 5x - 44


Hence we can say that,

2x + 1 = 5x - 44

3x = 45

x = 15


Thus PQ = 2*(15) + 1 = 31

PQ = 31